A 0.1 molar solution of NaCl is isotonic with 1% urea solution. To find out the van't Hoff factor and degree of ionisation of NaCl.


1
Expert's answer
2021-01-21T03:48:04-0500

NaCl(1−x) dissociates into Na+(x) and Cl(x)

Total concentration in solution =c(1+x) =0.1M(1+x)

Osmotic pressure of NaCl solution =0.1M(1+x) RT

Osmotic pressure of urea solution =0.1L (1/60) mol​RT

since solutions are isotonic, therefore, 1+x=1.667

(a) Van’s Hoff factor= 1.667

(b) Degree of dissociation =x=0.667